Assume youve generated the following information about the stock of Bens Banana Splits: The companys latest dividends of $Only part e please
Answer #1

PV of Price =24.72
FV or Price after 3 years =55.10
Dividend year 1 (D1) =1.80
D2 =1.93
D3 =2.07
Total cash flow in year 3 =2.07+55.10 =57.17
Using Financial calculator to calculate rate of return
CF0 =-24.72; CF1=1.80;CF2=1.93;CF=57.17;CPT IRR =36.72%

Rate of return =36.72%

If holding period return is required then =(55.10-24.72+1.80+1.93+2.07)/24.72 =146.36%